Make New Entry,
Make Followup Entry
User name Armstrong
Log entry time 22:42:51 on October 24, 2009
Entry number 296504
keyword=Left Q2 trips, but at higher current
OK, we made some improvement - we got up to 1125 A before Left Q2 trips,
with the new lead flows. This is a big improvement, since we had before
had it tripping at 668 A, so we have almost doubled the current by
increasing the lead flows from 95-100 l/m to 110l/m.
So, this gives us some hope, but we decide not to go any further, but to
establish beam on the Loop 1 target...